OpenStudy (anonymous):

@amistre64 @Luigi0210 @mathstudent55 @rrp240 aliyah works for isurvey.com she earned 35 points shes wants to refer people so she can get 2300 points if each referral is 20 points how much people will she need to refer to get to 2300 points

OpenStudy (anonymous):

114

OpenStudy (anonymous):

how

OpenStudy (austinl):

2x+35=2300 Solve for x.

OpenStudy (austinl):

20x*

OpenStudy (austinl):

\(20x+35=2300\) \(20x=2265\) \(x=\dfrac{453}{4}\approx113.25\) Because you cannot have \(.25\) people, you round up, 114 people.